# Towards Neural Synthesis for SMT-Assisted Proof-Oriented Programming

### v1

@@ Abstract
Seeking to spur research on using AI to automate the construction of proof-oriented programs, we curate a dataset of 600K lines of open-source F⋆ programs and proofs, including software used in production systems ranging from Windows and Linux, to Python and Firefox. Our dataset includes around 32K top-level F⋆ definitions, each representing a type-directed program and proof synthesis problem—producing a definition given a formal specification expressed as an F⋆ type. We provide a program-fragment checker that queries F⋆ to check the correctness of candidate solutions. We believe this is the largest corpus of SMT-assisted program proofs coupled with a reproducible program-fragment checker.
@@
1. A new dataset of programs and proofs: Aiming to spur research in AI-assisted proof-oriented programming, our first main contribution is FStarDataSet, a dataset of F⋆ programs and proofs extracted from 2060 source files, representing about 600K lines of source code, drawn from 8 open source projects. The dataset provides around 32K top-level F⋆ definitions, coupled with tooling that allows each definition to be checked in isolation. We believe this is the largest dataset of SMT-assisted proof-oriented programs and we envision for it to be a live, evolving data set, with new projects added to it over time. Indeed, even in the period of preparing this paper, FStarDataSet has grown to include 4 additional projects reaching 940K lines of source code. Although we currently focus on F⋆, we hope for the dataset to also grow to include data from other proof-oriented languages. We describe the dataset in detail in §III.
@@
Based on our results, we are optimistic about for the future of AI-assisted proof-oriented programming. Researchers building verified systems in proof-oriented languages have reported writing around 3-5 lines of proof for each line of verified code [24, 25], a considerable overhead, despite strong symbolic automation from SMT solvers. For SMT-assisted proof-oriented programs, we provide the first, substantial empirical evidence that LLMs trained on corpora like FStarDataSet, and prompted with retrieval augmentation, can automate somewhere between a third and a half of proofs. That said, our approach focuses on synthesizing program and proof fragments given their specifications: finding the right modular decomposition to prove a program correct, with the right specifications and auxiliary lemmas, is not yet within the scope of the techniques we explore. §VI provides further discussion and analysis. We anonymously release the source code, scripts, and data for the experiments in this paper in https://doi.org/10.5281/zenodo.10853636.
@@ IV Neural Program & Proof Synthesis for F⋆
Prompt Preparation. We prepare the input to the LLM (colloquially known as a prompt), by combining the “goal” type, the file context, selected premises, retrieved related examples, together with some fixed instructions for the model. We reserve a fixed number of tokens in the prompt for the file context, premises, and related examples. The file context contains opened and instantiated modules in the file together with other definitions in the file up to the point of the goal type, preferring definitions that are closer to the goal to fit in the fixed token budget for the file context. Likewise, for related examples and premises, we rank them according to descending similarity values, including as many as fit into their respective token budgets. We organize different components within a natural language prompt, i.e., each component of the prompt is preceded by a brief description of what the component is. Finally, the prompt ends with a set of instructions including the prefix of the definition. Figure 1 shows an example prompt at bottom right. We experiment on the impact of different components (context, retrieved examples, and selected premises) in §V-C.

### v3

@@ Abstract
Seeking to spur research on using AI to automate the construction of proof-oriented programs, we curate a dataset of 600K lines of open-source F⋆ programs and proofs, including software used in production systems ranging from Windows and Linux, to Python and Firefox. Our dataset includes around 32K top-level F⋆ definitions, each representing a type-directed program and proof synthesis problem—producing a definition given a formal specification expressed as an F⋆ type. We provide a program-fragment checker that queries F⋆ to check the correctness of candidate solutions. We also report on an extended version of our dataset containing a total of 940K lines of programs and proofs, with a total of 54k top-level F⋆ definitions. We believe this is the largest corpus of SMT-assisted program proofs coupled with a reproducible program-fragment checker.
@@
1. A new dataset of programs and proofs: Aiming to spur research in AI-assisted proof-oriented programming, our first main contribution is FStarDataSet, a dataset of F⋆ programs and proofs extracted from 2060 source files, representing about 600K lines of source code, drawn from 8 open source projects. The dataset provides around 32K top-level F⋆ definitions, coupled with tooling that allows each definition to be checked in isolation. We believe this is the largest dataset of SMT-assisted proof-oriented programs and we envision for it to be a live, evolving data set, with new projects added to it over time. Indeed, currently FStarDataSet has grown to include 4 additional projects reaching 940K lines of source code (see Section -A, and -B for details and initial results). Although we currently focus on F⋆, we hope for the dataset to also grow to include data from other proof-oriented languages. We describe the dataset in detail in §III.
@@
Based on our results, we are optimistic about for the future of AI-assisted proof-oriented programming. Researchers building verified systems in proof-oriented languages have reported writing around 3-5 lines of proof for each line of verified code [24, 25], a considerable overhead, despite strong symbolic automation from SMT solvers. For SMT-assisted proof-oriented programs, we provide the first, substantial empirical evidence that LLMs trained on corpora like FStarDataSet, and prompted with retrieval augmentation, can automate somewhere between a third and a half of proofs. That said, our approach focuses on synthesizing program and proof fragments given their specifications: finding the right modular decomposition to prove a program correct, with the right specifications and auxiliary lemmas, is not yet within the scope of the techniques we explore. §VI provides further discussion and analysis. We release FStarDataSet in https://huggingface.co/datasets/microsoft/FStarDataSet. Details of the project, source code and models can be found in http://fstar-lang.org/popai
@@ IV Neural Program & Proof Synthesis for F⋆
Prompt Preparation. We prepare the input to the LLM (colloquially known as a prompt), by combining the “goal” type, the file context, selected premises, retrieved related examples, together with some fixed instructions for the model. We reserve a fixed number of tokens in the prompt for the file context, premises, and related examples. The file context contains opened and instantiated modules in the file together with other definitions in the file up to the point of the goal type, preferring definitions that are closer to the goal to fit in the fixed token budget for the file context. Likewise, for related examples and premises, we rank them according to descending similarity values, including as many as fit into their respective token budgets. We organize different components within a natural language prompt, i.e., each component of the prompt is preceded by a brief description of what the component is. We represent each of the related example with its type and definition. For each of the selected premises, we only use the short name of that premise in the prompt. While the type and definition of a premise contain more important information about it, we experimentally observe that using other information actually hurt the ultimate performance. Our conjecture is that since the number of allowed tokens for the prompt is limited, including additional information with each premise allows fewer premises to be included in the prompt, which leads to important premises being excluded from the prompt altogether. Finally, the prompt ends with a set of instructions including the prefix of the definition. Figure 1 shows an example prompt at bottom right. We experiment on the impact of different components (context, retrieved examples, and selected premises) in §V-C.
@@ VI-C Definition ingredients in the prompt and its impact
VI-E Alternative Prompt Preparation
[…]
Structured | 30.76 | 19.45 |
[…]
In addition to presenting a natural language prompt (NL prompt) (see §IV for details), we investigated alternative prompting strategies. In particular, we created a structured format format similar to the format proposed by Gupta et al. [48]. In this version, every input modalities are surrounded by tags such as <context> and </context>, surrounding the file context <related> and </related> around the related examples. In another version, we create a completion style where all the prompt components are concatenated with a separator token <|end_of_text|>, the default padding token of Phi-2, without any description of each of the components. Natural language prompts are slightly better that structured prompts, while the completion-style prompt performs substantially worse. We conjecture that since the Phi-2 model is pre-trained mostly on natural language, wrapping the prompt components with natural language help during fine-tuning; the structured tag formats also help the model distinguish the components. Such a distinction is lost in the completion style prompt.